Please help me solve
(4x)^2 + sqrt(2x+3) = 8x + 1
Thanks a lot, i'm really appreciate your help!

Answer by stanbon(75887)

(4x)^2 + sqrt(2x+3) = 8x + 1
----
sqrt(2x+3) = -4x^2+8x+1
---
2x+3 = (-4x^2+8x+1)(-4x^2+8x+1)
-----
2x + 3 = 16x^4-32x^3-4x^2 -32x^3+64x^2 + 8x - 4x^2+8x+1
2x+3 = 16x^4 - 64x^3 + 56x^2 + 16x + 1
----
16x^4 - 64x^3 + 56x^2 + 14x -2 = 0
----
8x^4 - 32x^3 + 28x^2 + 7x - 1 = 0
--------------
I graphed it and found zeroes at
x = -0.28.., x = +0.104.., x = 1.78.. and x = 2.40...
------
Check for extraneous zeroes.
Cheers,
Stan H.
================